OpenStudy (anonymous):

Identify the system of equations used to solve the following problem. How much of a 60% saline solution would you need to add to pure water to create 10 L of a 50% saline solution?

OpenStudy (mathstudent55):

The first step is to choose variables. You need to use some amount of 60% saline solution and some amount of water. Let the amount of saline solution be x, and let the amount of water be y. You are creating 10 L of solution. Using x and y as the amounts you will be mixing, and 10 as the end result, can you come up with an equation for the amounts of solutions and water?

OpenStudy (mathstudent55):

You will use x amount of 60% solution and y amount of water. Altogether, you will make 10 L of 50% solution. The first equation is: x + y = 10 Do you understand it so far?

OpenStudy (anonymous):

yes, im just terrible with %'s

OpenStudy (mathstudent55):

Ok. The first equation just dealt with the amounts of the 60% solution and water needed to produce 10 L of 50% solution. Now we need to deal with the salt content. First, look at the final result. 10 L of 50% solution contains 50% * 10L = 0.50 * 10L = 5 L of salt.

OpenStudy (mathstudent55):

The final solution will will be 10 L of 50% solution which contains 5 L of salt. The first solution is 60% solution. We will use x liters of it. How much salt is there in x liters of 60% solution? What is 60% of x written as an algebraic expression using x?

OpenStudy (mathstudent55):

60% of x is 60% * x = 0.6 * x = 0.6x The amount of salt in x liters of 60% solution is 0.6x. Water contributes no slat, so we get 0% * y = 0 The total salt in the end solution is 5 L Now we have all we need to write the second equation 0.6x + 0y = 5 or simply 0.6x = 5 That means the system of equations is: x + y = 10 0.6x = 5